# Finding Nontrivial Minimum Fixed Points in Discrete Dynamical Systems

### 08 Dec 2023 v3

@@ Abstract
Abstract. Networked discrete dynamical systems are often used to model the spread of contagions and decision-making by agents in coordination games. Fixed points of such dynamical systems represent configurations to which the system converges. In the dissemination of undesirable contagions (such as rumors and misinformation), convergence to fixed points with a small number of affected nodes is a desirable goal. Motivated by such considerations, we formulate a novel optimization problem of finding a nontrivial fixed point of the system with the minimum number of affected nodes. We establish that, unless P = NP, there is no polynomial time algorithm for approximating a solution to this problem to within the factor $n^{1-\epsilon}$ for any constant $\epsilon>0$. To cope with this computational intractability, we identify several special cases for which the problem can be solved efficiently. Further, we introduce an integer linear program to address the problem for networks of reasonable sizes. For solving the problem on larger networks, we propose a general heuristic framework along with greedy selection methods. Extensive experimental results on real-world networks demonstrate the effectiveness of the proposed heuristics.
@@ 1 Introduction
Emails: {zq5au, vsakumar,marathe}@virginia.edu, {chenannie45, drosenkrantz, ssravi0, thestearns2}@gmail.com
@@ 2 Related work
Fixed points of discrete dynamical systems have been widely studied. Goles and Martinez ([15]) show that that for any initial configuration, a threshold SyDS always converges to either a fixed point or a cycle with two configurations in a polynomial number of time steps. Barrett et al.([4]) show that determining whether a system has a fixed point (FPE) is NP-complete for symmetric sequential dynamical systems and that the problem is efficiently solvable for threshold sequential dynamical systems. More recently, Chistikov et al. ([11]) study fixed points in the context of opinion diffusion; they show that determining whether a system reaches a fixed point from a given configuration is PSPACE-complete for SyDSs on general directed networks, but can be solved in polynomial time when the underlying graph is a DAG. In Rosenkrantz et al. ([25]), we investigate convergence and other problems for SyDSs whose underlying graphs are DAGs. In particular, we show that the convergence guarantee problem (i.e., determining if a system reaches a fixed point starting from any configuration) is Co-NP-complete for SyDSs on DAGs.
@@
Existing works on influence minimization focus on reducing the prevalence via control strategies. Yang, Li and Giua ([31]) study the problem of finding a $k$-subset of active nodes (i.e., an initial configuration with at most $k$ nodes in state 1) such that the converged influence value is minimized and a target set of nodes are active. They provide an integer program of the problem and suggest two heuristics. Wang et al. ([28]) propose a new rumor diffusion model and optimize blocking the contagion by considering an Ising model. Zhu, Ni, and Wang ([35]) estimate the influence of nodes and minimize the adoption of negative contagions by disabling nodes. Other approaches focus on blocking the spread via node removal [21, 32, 8, 22] or edge removal [20, 19, 7, 23] and enhance network resilience [8].
@@
Agent decision-making in coordination games coincides with threshold-based cascade of contagions. Adam et al. ([1]) study the best response dynamics of coordination games and analyze the convergences and propose a new network resilience measure. Ramazi et al. ([24]) study both coordination and anticoordination games and show that such games always reach equilibria in a finite amount of time. Other aspects of equilibria in networked games (such as developing control strategies and determining the existence of equilibria) have also been studied ([34, 6, 2, 26]).
@@ 4 Computational Hardness of NMIN-FPE
This completes the construction of the NMin-FPE instance $\mathcal{S}=(G_{\mathcal{S}},\mathcal{F})$ which clearly takes polynomial time w.r.t. $n_{\mathcal{M}}$. The resulting graph $G_{\mathcal{S}}{}$ has $n_{\mathcal{S}}=n_{\mathcal{M}}+m_{\mathcal{M}}+\beta+2$ vertices and $m_{\mathcal{S}}=n_{\mathcal{M}}+3m_{\mathcal{M}}+\beta$ edges. Furthermore, $G_{\mathcal{S}}{}$ is bipartite. An example of reduction is shown in Figure 2.
We now argue that $G_{\mathcal{M}}$ has a vertex cover of size at most $k$ if and only if the algorithm $\mathcal{A}$ returns a non-zero fixed point of size at most $n_{\mathcal{S}}^{1-\epsilon}\alpha$ where $\alpha=m_{\mathcal{M}}+n_{\mathcal{M}}+1$.
@@ 6               $C^{*}=C_{u}$
The fixed point condition. We use a superscript to denote the iteration number. Initially, $A_{u}^{(0)}$ contains only $u$, and we actively add a new node to $A_{u}^{(k)}$ in each iteration $k\geq 1$. To determine if the fixed point condition is met at the $k$th iteration, we maintain a value $\delta^{(k)}$ that is the number of additional nodes that need to be selected to satisfy the thresholds of nodes in $A_{u}^{(k)}$ at the $k$th iteration. Heuristically, $\delta^{(k)}=\sum_{v\in A_{u}^{(k)}}\tilde{\tau}_{v}^{(k)}$ where $\tilde{\tau}_{v}^{(k)}=\max\{0,\tau_{v}-|A_{u}^{(k)}\cap N(v)|\}$ is the residual threshold of $v$ at the $k$th iteration. We can view $\tilde{\tau}_{v}^{(k)}$ as the additional number of $v$’s neighbors that need to be selected to satisfy $\tau_{v}$.
@@ 6.2 Experimental results
We remark that the heuristics achieve high efficiency by the pruning technique. For some networks, Gurobi runs comparably fast, usually within $5$ minutes. Nevertheless, Gurobi uses parallelization mechanisms such as multithreading (over $30$ threads are used on each instance) whereas our heuristics achieve high efficiency while executing in serial mode.

### 18 Feb 2024 v4

@@ Abstract
Networked discrete dynamical systems are often used to model the spread of contagions and decision-making by agents in coordination games. Fixed points of such dynamical systems represent configurations to which the system converges. In the dissemination of undesirable contagions (such as rumors and misinformation), convergence to fixed points with a small number of affected nodes is a desirable goal. Motivated by such considerations, we formulate a novel optimization problem of finding a nontrivial fixed point of the system with the minimum number of affected nodes. We establish that, unless P = NP, there is no polynomial time algorithm for approximating a solution to this problem to within the factor $n^{1-\epsilon}$ for any constant $\epsilon>0$. To cope with this computational intractability, we identify several special cases for which the problem can be solved efficiently. Further, we introduce an integer linear program to address the problem for networks of reasonable sizes. For solving the problem on larger networks, we propose a general heuristic framework along with greedy selection methods. Extensive experimental results on real-world networks demonstrate the effectiveness of the proposed heuristics.
Conference version. The conference version of the paper is accepted at AAAI-2022: link.
@@ 2 Related work
Fixed points of discrete dynamical systems have been widely studied. Goles and Martinez [15] show that for any initial configuration, a threshold SyDS always converges to either a fixed point or a cycle with two configurations in a polynomial number of time steps. Barrett et al. [4] show that determining whether a system has a fixed point (FPE) is NP-complete for symmetric sequential dynamical systems and that the problem is efficiently solvable for threshold sequential dynamical systems. More recently, Chistikov et al. [11] study fixed points in the context of opinion diffusion; they show that determining whether a system reaches a fixed point from a given configuration is PSPACE-complete for SyDSs on general directed networks, but can be solved in polynomial time when the underlying graph is a DAG. In Rosenkrantz et al. [25], we investigate convergence and other problems for SyDSs whose underlying graphs are DAGs. In particular, we show that the convergence guarantee problem (i.e., determining if a system reaches a fixed point starting from any configuration) is Co-NP-complete for SyDSs on DAGs.
@@
Existing works on influence minimization focus on reducing the prevalence via control strategies. Yang, Li and Giua [31] study the problem of finding a $k$-subset of active nodes (i.e., an initial configuration with at most $k$ nodes in state 1) such that the converged influence value is minimized and a target set of nodes are active. They provide an integer program of the problem and suggest two heuristics. Wang et al. [28] propose a new rumor diffusion model and optimize blocking the contagion by considering an Ising model. Zhu, Ni, and Wang [35] estimate the influence of nodes and minimize the adoption of negative contagions by disabling nodes. Other approaches focus on blocking the spread via node removal [21, 32, 8, 22] or edge removal [20, 19, 7, 23] and enhance network resilience [8].
@@
Agent decision-making in coordination games coincides with threshold-based cascade of contagions. Adam et al. [1] study the best response dynamics of coordination games and analyze the convergences and propose a new network resilience measure. Ramazi et al. ([24]) study both coordination and anticoordination games and show that such games always reach equilibria in a finite amount of time. Other aspects of equilibria in networked games (such as developing control strategies and determining the existence of equilibria) have also been studied [34, 6, 2, 26].
@@ 4 Computational Hardness of NMIN-FPE
This completes the construction of the NMin-FPE instance $\mathcal{S}=(G_{\mathcal{S}},\mathcal{F})$ which clearly takes polynomial time w.r.t. $n_{\mathcal{M}}$. The resulting graph $G_{\mathcal{S}}{}$ has $n_{\mathcal{S}}=n_{\mathcal{M}}+m_{\mathcal{M}}+\beta+2$ vertices and $m_{\mathcal{S}}=n_{\mathcal{M}}+3m_{\mathcal{M}}+\beta$ edges. Furthermore, $G_{\mathcal{S}}{}$ is bipartite. An example of reduction is shown in Figure 2. We now argue that $G_{\mathcal{M}}$ has a vertex cover of size at most $k$ if and only if the algorithm $\mathcal{A}$ returns a non-zero fixed point of size at most $n_{\mathcal{S}}^{1-\epsilon}\alpha$ where $\alpha=m_{\mathcal{M}}+n_{\mathcal{M}}+1$.
@@ 6               $C^{*}=C_{u}$
The fixed point condition.
We use a superscript to denote the iteration number. Initially, $A_{u}^{(0)}$ contains only $u$, and we actively add a new node to $A_{u}^{(k)}$ in each iteration $k\geq 1$. To determine if the fixed point condition is met at the $k$th iteration, we maintain a value $\delta^{(k)}$ that is the number of additional nodes that need to be selected to satisfy the thresholds of nodes in $A_{u}^{(k)}$ at the $k$th iteration. Heuristically, $\delta^{(k)}=\sum_{v\in A_{u}^{(k)}}\tilde{\tau}_{v}^{(k)}$ where $\tilde{\tau}_{v}^{(k)}=\max\{0,\tau_{v}-|A_{u}^{(k)}\cap N(v)|\}$ is the residual threshold of $v$ at the $k$th iteration. We can view $\tilde{\tau}_{v}^{(k)}$ as the additional number of $v$’s neighbors that need to be selected to satisfy $\tau_{v}$.
@@ 6.2 Experimental results
The heuristics achieve high efficiency by the pruning technique. For some networks, Gurobi runs comparably fast, usually within $5$ minutes. Nevertheless, Gurobi uses parallelization mechanisms such as multithreading (over $30$ threads are used on each instance) whereas our heuristics achieve high efficiency while executing in serial mode.